Bongiorno’s is a pizzeria operating in a competitive market. Its short-run production function is shown below. Bongiorno’s weekly cost of capital (plant) is $750, its cost of labor is $320 per worker, and its cost of the pizza ingredients is $6 per pie. The going market price of pizza is $13 per pie.

Workers (L) Pizzas (Q) 132 200 255 303 347 387 424

A) In the scenario above, when Bongiorno’s hires a fourth worker, what is the marginal product of labor?

B) In the scenario above, if Bongiorno’s hires a fifth worker, what will be the value of the marginal product of labor?

C) In the scenario above, how many workers should Bongiorno’s hire?

D) In the scenario above, when Bongiorno’s produces 255 pizzas, what is its total variable cost ($)?

E) In the scenario above, when Bongiorno’s produces 347 pizzas, what is its average total cost ($)?

F) In the scenario above, suppose Bongiorno’s temporarily shuts down and does not produce any pizza. What is its total cost ($)?

G) In the scenario above, what price should Bongiorno’s charge to maximize its profit ($ per pizza)?

H) In the scenario above, which of the following output levels will bring Bongiorno’s the highest profit?

(a) 132 pizzas (b) 200 pizzas (c) 255 pizzas (d) 303 pizzas (e) 347 pizzas (f) 387 pizzas

I) In the scenario above, what is the highest profit Bongiorno’s can achieve ($)?
Answer #1
Workers Pizzas MP VMPL TVC TFC TC ATC
1 132 132 1716 1112 750 1862 14.11
2 200 68 884 1840 750 2590 12.95
3 255 55 715 2490 750 3240 12.71
4 303 48 624 3098 750 3848 12.70
5 347 44 572 3682 750 4432 12.77
6 387 40 520 4242 750 4992 12.90
7 424 37 481 4784 750 5534 13.05

MP = change in output

VMPL = MP*P

TVC=(320*Workers)+(6*Pizzas)

TFC=750

TC=TVC+TFC

ATC=TC/Q

A) MP = 48

B) VMPL = 572

C) It should hire where VMPL is equal to or greater than the wage of the worker which is 320 so it should hire 7 workers

D) TVC = 2490

E) ATC = 12.77

F) TC = fixed cost = 750

As per HOMEWORKLIB RULES, more than four parts are answered